OpenStudy (anonymous):

Find the point(s) of intersection: x^2 + y^2 = 25 xy = 12

OpenStudy (tkhunny):

Have you considered substitution? From the second, x = 12/y -- Go!

OpenStudy (anonymous):

I tried, but I do not know how to evaluate 144/y^2 + y^2 = 25

OpenStudy (anonymous):

\[\frac{y^4+144}{y^2}=25\] \[y^4+144=25y^2\] \[y^4-25y^2+144=0\]

OpenStudy (tkhunny):

What did you try? Can you solve 144/x+ x = 25?

OpenStudy (anonymous):

From y^4 - 25^2 + 144 = 0 would you factor? If so, how?